|10/02/2002||Services on the Gold server were unavailable from 4:40 AM until 4:59 AM this morning due to system maintenance.|
|09/27/2002||Intermittent problems have been reported with some Truman computer systems not synchronizing their time with the Truman time server. We are investigating the problem. ** Updated 9-25 ** The time sync problem has been identified and resolved.|
|09/27/2002||The slow internet access that some users experienced on Wednesday afternoon has been resolved. The cause was a "Denial-of-Service" attack originating from a computer located on the Truman Network. The system has been disconnected while the issue is being investigated.|
|09/25/2002||All internal network access was disrupted on Monday, 9/23 for approximately 20 minutes at 3:00PM. The cause is currently being investigated, but looks to be due to fautly network hardware located in the Data Center.|
|09/26/2002||Internet access was down from 5:14PM until 5:16PM on 9/24/2002 to address the issue of slow response times when accessing certain Internet-based resources. Access was immediately restored and response times were back within normal ranges.|
|09/23/2002||Messages sent to Campus Pipeline users were delayed and a small number of messages may have bounced to the original sender beginning on Thursday evening, September 19. These bounced messages and all queued messages should be delivered to their intended recipient. The issue was resolved with a correction of the DNS configuration.|
|09/23/2002||Due to a system error on one of the SQL servers, selected dynamic websites - including ITS - were not funtional from approximately 7:30 to 8:30 Friday morning.|
|09/18/2002||CenturyTel, a MOREnet provisioned frame relay provider in the Springfield area, is currently experiencing a fiber cut. Most MOREnet customers in the Springfield area have either lost service or can expect severe network latency during this circuit problem. MOREnet is currently working with the provisioner to resolve this matter as quickly as possible.|
|09/20/2002||Some Network services will be intermittently disrupted from 6:00PM until 12:00AM on Thursday, September 19, 2002. Upgrades and normal maintenance will be done on several network servers including the DNS, DHCP and WINS Servers. Other maintenance will include upgrades to the Data Center Network.
During this period, users may see slow response times when accessing the Truman network and Internet and/or may not be able to access network or Internet resources.
All network services should be restored after midnight on Thursday night.
|09/14/2002||Due to a critical network hardware failure, Internet access was temporarily disrupted this afternoon from 2:30 until 3:00. All service has since been restored.|
|09/11/2002||Due to problems with the electronic printer accounting system, page charges for student printing has been temporarily suspended.|
|09/12/2002||Campus Pipeline will be down this afternoon from 2:00PM until 3:00PM for critical maintenance.
Campus Pipeline, Inc has supplied a patch that is supposed to fix the Calendar and the Group problems. |
|09/05/2002||There continues to be problems with the Groups section of Campus Pipeline. After a period of time, the Groups section is no longer available to users. Campus Pipeline, Inc. continues to work on this issue, but has yet to identify the problem or release a patch. Updates will be here, and in the announcement section of pipeline when the problem is resolved.|
|09/03/2002||MOREnet network circuits are currently down in St. Louis. Internet sites in St. Louis or Internet traffic going through St. Louis may not work. MOREnet is working on resolving the problem as soon as possible. ** Update ** at ~ 2:45 pm MOREnet notified us that they have fixed the problem.|
|09/04/2002||Gold was down for maintenance for approximately 45 minutes between 5am and 6am on Tuesday 9.3.02. Faculty/Staff e-mail and all personal websites were unavailable during the maintenance window.|
|09/10/2002||We have experienced random intermittent problems with slow Internet access. We are working on the problem and have opened a ticket with MOREnet. MOREnet is aware of their problems with the state network. If your system is affected, you may be able to use the Truman Proxy Server to continue accessing the Internet. Instructions for using the Truman Proxy Server are located at: http://its.truman.edu/documentation/remote_computing/remote_library.stm |
|09/08/2000||The network print server for the trumannt domain was down yesterday (9/5) for an indeterminate amount of time. Service was restored at 3:05pm.|
|09/08/2000||There has been a critical hardware failure in the mail server, Tin. This has brought all incoming and outgoing mail to a halt. Truman mail set to be forwarded to an internet account cannot be delivered. IBM is on-site, hoping to repair or replace the equipment as soon as possible.|
|09/08/2000||IBM techs have determined that Tin's failure was due to a bad motherboard. Additional IBM techs should arrive with the new board by 1:00 today. It is hoped that mail service will be restored by 2:00pm.|
|09/08/2000||Truman ITS staff have moved mail services to an alternate server and have fully restored e-mail services!! Any addresses in your mail address books which include Tin (e.g. "email@example.com") will not function. Those addresses should be modified to the standard "firstname.lastname@example.org". |
|09/10/2000||he Mac server, Neon, experienced problems from ~2:00pm
Wednesday (9/6) till 4:00pm Thursday (9/7). These problems
prevented log-ons in academic and res hall labs. The problems are
being addressed, but may require some planned downtime in the near
future to fully fix. Those downtimes will be publicized and are planned
for low-usage time periods. |
|09/10/2000||Several areas of Barnett Hall network were down from 8:00am to
10:15am. A contractor had unplugged a network electronics
component from electrical power.|
|09/14/2000||The network was down in Fair apartments Monday night until Tuesday morning. Electrical power was cut to the wiring closet. The power strip was shut off and removed from the wiring cabinet by an unknown party. |
|09/15/2001||The network was down in Pickler Library first floor, east wing today for
several hours due to an electrical circuit overload. The circuit breaker
was reset at ~ 10:30 AM.|
|09/15/2000||The Mac server, Neon, is currently down for repair and maintenance.
There was an unexpected problem regarding disk space which occurred Tuesday
evening (9.12.2000) ITS technicians are working to bring this server|